Two men are lifting a weight M of 1,000 kg at a co
Answer #1

here,

mass of the block , m = 1000 kg

at time instant t1

equating the foces horizontally

F1 *cos(80) - F2 *cos(60) = 0 ...(1)

equating the forces vertically

F1*sin(80) + F2*sin(60) - 1000*9.8 = 0 ...(2)

from equation (1) and (2)

F1 = 7623.05 N

F2 = 2647.46 N

at time t2

the angle of F1 with the horizontal , theta1 = arctan((56.71 - 10)/(56.71/tan(80)))

theta 1 = 77.92 degree

the angle of F2 with the horizontal , theta2 = arctan((17.32 - 10)/(17.32/tan(60)))

theta2 = 36.2 degree

equating the foces horizontally

F1 *cos(77.92) - F2 *cos(36.2) = 0 ...(1)

equating the forces vertically

F1*sin(77.92) + F2*sin(36.2) - 1000*9.8 = 0 ...(2)

from equation (1) and (2)

F1 = 8664.71 N

F2 = 2247.11 N